P&M Inc. 2012 Income Statement Net Sales 36,408 Cost of Goods Sold 28,225 Depreciation 1,760 EBIT 6,423 Interest Expense 510 EBT 5,913 Taxes 2,070 Net Income 3,843 P&M Inc. 2012 Balance Sheet Cash 1,003 Account Payable 3,668 Account Receivable 4,218 Accrued 4,716 Inventory 21,908 Current Liability 8,384 Current Assets 27,129 Long-term Debt 11,500 Net Fixed Assets 14,080 Common Stock 17,500 Retained Earnings 3,825 Total Assets 41,209 Total Liabilities and Equity 41,209 What is the equity to asset ratio for 2012? 0.5988 0.5814 0.5319 0.5181 0.4926
Answer #1

EQUITY TO ASSET RATIO = EQUITY/TOTAL ASSETS

EQUITY = COMMON STOCK + RETAINED EARNINGS = 17500 + 3825 = 21325

TOTAL ASSETS = 41209

EQUITY TO ASSET RATIO = 21325/41209 = 51.75%

ANSWER : 0.5181 (THUMBS UP PLEASE)